the top of a building is a rectangle whose length is 60ft.more than the width.the perimeter 520 ft. find the length and the width
Answer by ikleyn(52876)

.

W + (W+60) = 520/2


2W + 60 = 260


2W = 260 - 60


2W = 200


W = 100.


ANSWER.  The width is 100 ft;  the length is  100+60 = 160 ft.